OpenStudy (anonymous):

Simplify: (2x2 – 6x – 20)/( 2x2 – 4x – 30) Select one: 2(x + 3) 2(x – 5)/(x + 3) (x + 2)/(x + 3) (x + 2)/(x – 5)

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{2x^2 – 6x – 20}{2x^2 – 4x – 30} } \] this?

OpenStudy (anonymous):

ok

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{2(x^2 – 3x – 10)}{2(x^2 – 2x – 15)} } \] \[\huge\color{blue}{ \frac{\cancel{2 } (x^2 – 3x – 10)}{\cancel{2 } (x^2 – 2x – 15)} } \] \[\huge\color{blue}{ \frac{ (x^2 – 3x – 10)}{ (x^2 – 2x – 15)} } \]so far so good?

OpenStudy (anonymous):

yeaf!

OpenStudy (solomonzelman):

\[\huge\color{blue}{ \frac{(x+2)(x-5)}{ (x+3)(x-5)} } \]I factored top and bottom, \[\huge\color{blue}{ \frac{(x+2)\cancel{(x-5)} }{ (x+3)\cancel{(x-5)} } } \]\[THE ~~~~ANSWER ~~~~IS?\]

OpenStudy (solomonzelman):

If you can't tell me what the answer is now, then you are not prepared and don't have enough knowledge to do this type of problems (sorry to say this) Or perhaps, you went to eat or something else, so then, I apologize, and take my words back, forgive me.

OpenStudy (anonymous):

(x + 2)/(x + 3) @SolomonZelman

OpenStudy (solomonzelman):

YES!!!!!!!!!!!!!!

OpenStudy (anonymous):

haha yeah thanks :)

OpenStudy (solomonzelman):

SO, the perhaps was right..

OpenStudy (anonymous):

I was working on other problems

OpenStudy (solomonzelman):

I see, You welcome, good work!